# Reconfigurable Intelligent Surface Aided Vehicular Edge Computing: Joint Phase-shift Optimization and Multi-User Power Allocation

### v1

@@ III-B Queue Model
For VU $k$, we assume that its task arrival rate at time slot $t$ is ${\eta_{k}}(t)$, which follows a Poisson distribution, so one can calculate the tasks of VU $k$ during time slot t as
${a_{k}}(t)={\eta_{k}}(t)\times\Delta t.$ (1)
The arriving tasks will be saved in the buffer, and processed at time slot $t+1$, thus the buffer length of VU $k$ at time slot $t+1$ is given by
@@
where ${q_{k,o}}(t)$ and ${q_{k,l}}(t)$ denote the amount of tasks processed by task offloading and locally, respectively. Note that ${\left[x\right]^{+}}=\max(0,x)$.
@@ III-C Offloading Model
In the proposed model, it is assumed that the location of the $k$th vehicle is $(x_{k}^{t},y_{k}^{t},z_{k}^{t})$ at time slot $t$, and the coordinates of the BS and the RIS are $(B{S_{x}},B{S_{y}},B{S_{z}})$ and $(RI{S_{x}},RI{S_{y}},RI{S_{z}})$, respectively. Therefore, the channel gain between the $\mathit{k}$th VU and BS at slot $\mathit{t}$ is
@@
where ${d_{r,b}}$ is the geometric distance from the RIS to the BS, ${\alpha_{r,b}}$ is the path loss exponent of the RIS-BS link, and $R$ is the Rician coefficient associated with small-scale fading. LoS components ${\bf{h}}_{r,b}^{LoS}$ is defined as follows:
@@
${\bf{h}}_{k,r}^{t}=\sqrt{\rho{{\left({d_{k,r}^{t}}\right)}^{-{\alpha_{k,r}}}}}\sqrt{\frac{R}{{1+R}}}{\bf{h}}_{k,r}^{t\;LoS},\forall k\in K,\forall t\in T.$ (6)
Here, $d_{k,r}^{t}$ is the geometric distance between the $k$th vehicle and the RIS at time slot $t$, and ${\alpha_{k,r}}$ is the path loss exponent between the vehicle and the RIS. Note that ${\bf{h}}_{k,r}^{t\;LoS}$ is expressed as follows
@@
In this paper, we consider direct communication links coexisting with RIS-assisted communication links. Thus, we can obtain the signal-to-noise ratio (SNR) between the $k$th VU and the BS at time slot $t$ as
@@
When VU $k$ chooses to offload tasks to the VEC server associated with the BS, based on the formula (7), we can obtain the number of offloaded tasks processed by VU $k$ at time slot $t$ as
@@ III-D Local Execution
When VU $k$ selects to process tasks locally, we can obtain the size of tasks that can be processed locally at time slot $t$ as
@@
where ${p_{k,l}}(t)\in[0,{P_{\max,l}}]$ is the local execution power of VU $k$ at time slot $t$, and $c$ is the effective selection capacitance.
@@ III-E Problem Formulation
The target is to minimize the power consumption of task offloading and local processing as well as buffer length. Therefore, the multi-objective optimization problem for each VU $k$ can be formulated as follows:
@@
where (12b) and (12c) represent the power constraints for VU $k$ when offloading tasks and processing tasks locally, respectively. Due to the limitation of RIS hardware, the size of RIS phase-shift can only be selected within a limited range constrained by (12d). Furthermore, the objective function is non-convex, so this optimization problem is difficult to be solved. To better address this issue, we propose a novel framework by combining the modified MADDPG algorithm with the centralized DDPG algorithm. Since the RIS phase-shift controller and VUs are not the same type of agents, their actions and reward are also different. Therefore, we will first use the DDPG algorithm to obtain the optimal coefficient of phase-shift, and then obtain the optimal power allocation scheme through the modified MADDPG algorithm.
@@ IV RIS Phase-shift Optimization Based On DDPG
RIS is used as a tool to assist vehicle communication, and its goal is to maximize the information transmission rate of the vehicle by constantly adjusting the phase-shift coefficients to improve the performance of the system. Firstly, we fix the offloading power of the VUs, and then with the objective of maximizing the information transmission rate of the VU communication link, we use the DDPG algorithm to train the phase-shift of the RIS, so as to obtain the optimal phase-shift coefficients of the RIS. Thus, the phase-shift coefficient optimization problem for RIS can be expressed as:
@@
$\quad{\rm{}}\theta_{n}^{t}\in\Phi,{\rm{}}\forall n\in\mathcal{N},{\rm{}}\forall t\in\mathcal{T},$ (13c)
@@ IV-A DDPG for RIS Phase-shift Coefficients
The BS acts as an agent interacting with the RIS-assisted vehicular network and adjusts the phase-shift coefficients of the RIS according to its strategy, aiming at solving the optimization problem P2. At time slot $t$, the BS obtains the current state $s_{t}$, and then obtains the corresponding action based on the current selected strategy, and transit to the next state $s_{t+1}$ and also obtains the reward $r_{t}$. this process can be formulated as ${e_{t}}=({s_{t}},{a_{t}},{r_{t}},{s_{t+1}})$. The relevant state space, action space, and reward function in this model are represented as follows:
State space: the state obtained by the BS at time slot $t$ consists of the following components: phase-shift coefficient of the RIS $\theta_{n}^{t}$ , planar coordinate position of the VU $k$ $(x_{k}^{t},y_{k}^{t})$. In addition, through equation (8), the SNR of VU $k$ at time slot $t$, i.e. ${\gamma_{k}}(t-1)$ , depends on ${h_{r,b}}$ and $h_{k,r}^{t}$, which reflects the channel uncertainty of the VU at time slot $t$. Therefore, the state space of the agent is ${s_{t}}=[\theta_{1}^{t},\cdots,\theta_{N}^{t},x_{1}^{t},y_{1}^{t},\cdots,x_{K}^{t},y_{K}^{t},{\gamma_{1}}(t-1),\cdots,{\gamma_{K}}(t-1)].$
@@
Initialize target networks by $\theta_{\mu^{*}}\leftarrow\theta_{\mu}$, $\theta_{Q^{*}}\leftarrow\theta_{Q}$;
@@ IV-B Concept of DDPG Algorithm
For high-dimensional state spaces, DDPG can be used to predict appropriate actions given the current state by deep neural networks that can learn complex features and patterns of the state. Therefore for the high-dimensional state and action space in this model, we use DDPG for training, and after obtaining the continuous action values, we can obtain the appropriate RIS phase-shift coefficients by scaling and rounding. In the DDPG algorithm, the state space and action space are usually represented by a deep neural network with two networks, actor and critical: the actor network is used to learn the strategy (i.e., action selection), and its goal is to maximize the estimation of the action value by the critical network, and the critical network is used to evaluate the value of the strategy, and learns to assess the evaluation by means of supervised learning merit of the actions output by the actor network. In addition, the DDPG algorithm uses a target actor network and a target critic network in order to ensure the stability of the algorithm. The network structure of these two networks is the same as that of the actor and critic networks, respectively.
Firstly, initialize the parameters $\theta_{\mu}$ and $\theta_{Q}$ of actor network $\mu(s;{\theta_{\mu}})$ and critical network $Q(s,a;{\theta_{Q}})$, and set the parameters $\theta_{\mu^{\prime}}$ and $\theta_{Q^{\prime}}$ of target actor network $\mu^{\prime}(s;\theta_{\mu^{\prime}})$ and target critic network $Q^{\prime}(s,a;\theta_{Q^{\prime}})$, respectively. Then, at each time slot $t$, action $a_{t}$ is selected based on the current strategy $\mu$ and exploration noise $\Delta n$, while the reward and next state are observed by executing the action. Next, the experience of each step is stored in the experience buffer for random sampling training. At each time slot $t$, when the buffer length is greater than $I$, a batch of data is sampled from the experience replay buffer $\mathcal{D}$ which is used to update the parameters of the critic network to minimize the loss function
@@
${\nabla_{{\theta_{\mu}}}}J\approx\frac{1}{{I}}\sum\limits_{i}{{\nabla_{a}}}Q(s,a;{\theta_{Q}}){|_{s={s_{i}},a=\mu({s_{i}})}}\nabla{\theta_{\mu}}\mu(s;{\theta_{\mu}}){|_{{s_{i}}}}$ (17)
@@ V Power Allocation Optimization By Modified MADDPG Algorithm
In this section, we will describe in detail the multi-agent environment and its associated states, actions and rewards, and finally, we will discuss the proposed MARL algorithm and its associated formulations. To be clear, our proposed framework for single-agent reinforcement learning and multi-agent reinforcement learning (SARL-MARL) is shown in Fig. 2.
@@ V-A Power Allocation Model for VEC
Algorithm 2 Modified MADDPG Algorithm for Power Allocation
In the previous section, the BS acted as an agent and interacted with the environment, whereas in this section, we take the VU as an agent and use a multi-agent approach to interact with the environment and obtain power allocation decisions through corresponding strategies. Similarly, the relevant state space, action space, and reward function in this model are represented as follows:
State space: At time slot $t$, the state of each VU $k$ (agent $k$) consists of the following components: buffer length $q_{k}(t)$, the size of offloaded executed tasks $q_{k,o}(t)$, the size of locally executed tasks $q_{k,l}(t)$, and the offloaded and locally processed task overflows ${{q_{k,o}}(t)+{q_{k,l}}(t)-{q_{k}}(t)}$, the SNR of VU $k$ at time slot $t$ ${\gamma_{k}}(t-1)$. Therefore, the state space of the VU $k$ at slot $t$ can be formulated as ${s_{k}(t)}=\left[{{q_{k}}(t),{q_{k,o}}(t),{q_{k,l}}(t),{q_{k,o}}(t)+{q_{k,l}}(t)-{q_{k}}(t),{\gamma_{k}}(t-1)}\right].$
Action space: as described above, agent $k$ allocates the offloading power and the local power according to the corresponding policy, so that the action space of agent $k$ at time slot $t$ is defined as ${a_{k}}(t)=\left[{{p_{k,o}}(t),{p_{k,l}}(t)}\right]$.
Reward function: in our proposed scheme, there are two aspects of rewards for each agent, one is the global reward reflecting the cooperation among agents, and the other is the local reward that is to help each agent to explore the optimal power allocation scheme. In DRL, rewards can be set flexibly, and a good reward can improve the performance of the system. Here, our goal is to optimize the agent $k$’s offloading and the local power level, as well as the corresponding buffer length. Thus the local and global rewards of VU $k$ at time slot $t$ are
@@ V-B Multi-agent Deep Reinforcement Learning Algorithms
Based on DDPG algorithm, we propose a modified MADDPG algorithm that learns how to maximize both global and local rewards. It contains two main critics, one is a global critic, which is shared among all the agents and inputs the states and actions of all the agents to maximize the global reward; the other is local critic, which receives the states and actions of the specific agents and evaluates the its rewards. In addition, we consider the impact of approximation errors in strategy and value updates on the global critic in MADDPG algorithm. Similar to [68], to improve the MADDPG algorithm, we employ twin-delay deterministic strategy gradient to replace the global critic.
Specifically, we consider a vehicular environment with $K$ vehicles (agents) and the policies for all agents are $\mathbf{\pi}=\{{\pi_{1}},{\pi_{2}},\cdots,{\pi_{K}}\}$. The agent $k$’s strategy $\pi_{k}$, Q-functions ${\mathop{\rm Q}\nolimits}_{{\phi_{k}}}^{k}$ and twin global critic Q-functions (${\mathop{\rm Q}\nolimits}_{\psi 1}^{{g_{1}}}$, ${\mathop{\rm Q}\nolimits}_{\psi 2}^{{g_{2}}}$) are parameterized by ${\theta_{k}}$, $\phi_{k}$, ${\psi_{1}}$ and $\psi_{2}$, respectively. For each agent, the modified policy gradient can be written as
@@
${y_{g}}={r_{g}}+\gamma{\mathop{\min}\limits_{j=1,2}}{\left.{Q_{\psi_{j}^{\prime}}^{{g_{j}}}({\bf{s^{\prime},a^{\prime}}})}\right|_{a_{k}^{\prime}=\pi_{k}^{\prime}(s_{k}^{\prime})}}.$ (24)
Here, ${\bf{\pi}}^{\prime}=\{\pi_{1}^{\prime},\pi_{2}^{\prime},\cdots,\pi_{K}^{\prime}\}$ is the target policy with parameter $\theta^{\prime}=\{\theta_{1}^{\prime},\theta_{2}^{\prime},\cdots,\theta_{K}^{\prime}\}$. Similarly, the local critic $Q_{{\phi_{k}}}^{k}$ of agent $k$ is updated to
${L^{k}}({\psi_{k}})={{\rm E}_{{s_{k}},{a_{k}},{r_{k}},s_{k}^{\prime}}}\left[{{{(Q_{{\phi_{k}}}^{k}({s_{k}},{a_{k}})-y_{l}^{k})}^{2}}}\right],$ (25)
@@
The detailed MARL algorithm is described in the table Algorithm 2.

### v2

@@ III-B Queue Model
For the $k$th VU, we assume that its task arrival rate at time slot $t$ is ${\eta_{k}}(t)$, which follows a Poisson distribution, so one can calculate the tasks of the $k$th VU during time slot $t$ as
${a_{k}}(t)={\eta_{k}}(t)\times\Delta t,$ (1)
the arriving tasks will be saved in the buffer, and processed at time slot $t+1$, thus the buffer length of the $k$th VU at time slot $t+1$ is given by
@@
where ${q_{k,o}}(t)$ and ${q_{k,l}}(t)$ denote the amount of tasks processed by task offloading and locally, respectively. Note that ${\left[x\right]^{+}}=\max(0,x)$. In addition, in the process of information transmission, the latency is related to the length of the buffer. For each time slot of buffer, if the processing of the remaining tasks is too long, it may cause a large backlog of data in other parts of the network, which in turn leads to an increase in the overall system latency.
@@ III-C Offloading Model
In the proposed model, it is assumed that the location of the $k$th VU is $(x_{k}^{t},y_{k}^{t},z_{k}^{t})$ at time slot $t$, and the coordinates of the BS and the RIS are $(B{S_{x}},B{S_{y}},B{S_{z}})$ and $(RI{S_{x}},RI{S_{y}},RI{S_{z}})$, respectively. Therefore, the channel gain between the $\mathit{k}$th VU and BS at slot $\mathit{t}$ is
@@
where ${d_{r,b}}$ is the geometric distance from the RIS to the BS, ${\alpha_{r,b}}$ is the path loss exponent of the RIS-BS link, and $R$ is the Rician coefficient associated with small-scale fading. LoS components ${\bf{h}}_{r,b}^{LoS}$ is defined as
@@
${\bf{h}}_{k,r}^{t}=\sqrt{\rho{{\left({d_{k,r}^{t}}\right)}^{-{\alpha_{k,r}}}}}\sqrt{\frac{R}{{1+R}}}{\bf{h}}_{k,r}^{t\;LoS},\forall k\in K,\forall t\in T,$ (6)
where, $d_{k,r}^{t}$ is the geometric distance between the $k$th vehicle and the RIS at time slot $t$, and ${\alpha_{k,r}}$ is the path loss exponent between the vehicle and the RIS. Note that ${\bf{h}}_{k,r}^{t\;LoS}$ is expressed as
@@
In this paper, multiple VUs in the VEC adopt the orthogonal frequency division multiplexing (OFDM) [66] technique to communicate with the BS, and thus the interference over different subbands is not considered in the communication model. In addition, we consider direct communication links coexisting with RIS-assisted communication links. Thus, we can obtain the signal-to-noise ratio (SNR) between the $k$th VU and the BS at time slot $t$ as
@@
When the $k$th VU chooses to offload tasks to the MEC server associated with the BS, based on the formula (7), we can obtain the number of offloaded tasks processed by the $k$th VU at time slot $t$ as
@@ III-D Local Execution
When the $k$th VU selects to process tasks locally, we can obtain the size of tasks that can be processed locally at time slot $t$ as
@@
where ${p_{k,l}}(t)\in[0,{P_{\max,l}}]$ is the local execution power of the $k$th VU at time slot $t$, and $c$ is the effective selection capacitance.
@@ III-E Problem Formulation
To improve overall system efficiency, our goal is to address multiple performance metrics simultaneously. Specifically, the objective is to minimize the power consumption associated with task offloading and local processing, as well as the buffer length. Therefore, the multi-objective optimization problem for each the $k$th VU can be formulated as follows:
@@
where (12b) and (12c) represent the power constraints for the $k$th VU when offloading tasks and processing tasks locally, respectively. Due to the limitation of RIS hardware, the size of RIS phase-shift can only be selected within a limited range constrained by (12d). Furthermore, the objective function is non-convex, so this optimization problem is difficult to be solved. To better address this issue, we propose a novel framework by combining the modified MADDPG algorithm with the centralized DDPG algorithm. We will first use the DDPG algorithm to obtain the optimal coefficient of phase-shift, and then obtain the optimal power allocation scheme through the modified MADDPG algorithm.
@@ IV RIS Phase-shift Optimization Based On DDPG
RIS is used as a tool to assist vehicle communication, and its goal is to maximize the information transmission rate of the vehicle by constantly adjusting the phase-shift coefficients to improve the performance of the system. In additional, P1 involves optimizing the vehicle’s offloading power, local execution power, and RIS phase-shift coefficients. Since the vehicle and RIS are different types of agents, so in order to solve P1, we first solve the RIS phase-shift optimal problem during vehicle offloading by fixing the vehicle offloading power and then with the objective of maximizing the information transmission rate of the VU offloading link, we use the DDPG algorithm to train the phase-shift of the RIS, so as to obtain the optimal phase-shift coefficients of the RIS. Thus, the phase-shift coefficient optimization problem for RIS can be expressed as:
@@
$\quad{\rm{}}\theta_{n}^{t}\in\Phi,{\rm{}}\forall n\in\mathcal{N},{\rm{}}\forall t\in\mathcal{T}.$ (13c)
@@ IV-A DDPG for RIS Phase-shift Coefficients
In training the RIS phase shift coefficients with the DDPG algorithm, the BS acts as an agent interacting with the RIS-assisted vehicular network and adjusts the phase-shift coefficients of the RIS according to its strategy, aiming at solving the optimization problem $\mathit{P}2$. At time slot $t$, the BS obtains the current state $s_{t}$, and then obtains the corresponding action based on the current selected strategy, and transit to the next state $s_{t+1}$ and also obtains the reward $r_{t}$. this process can be formulated as ${e_{t}}=({s_{t}},{a_{t}},{r_{t}},{s_{t+1}})$. The relevant state space, action space, and reward function in this model are represented as follows:
State space: the state obtained by the BS at time slot $t$ consists of the following components: phase-shift coefficient of the RIS $\theta_{n}^{t}$ , planar coordinate position of the $k$th VU $(x_{k}^{t},y_{k}^{t})$. In addition, through equation (8), the SNR of VU $k$ at time slot $t$, i.e. ${\gamma_{k}}(t-1)$ , depends on ${h_{r,b}}$ and $h_{k,r}^{t}$, which reflects the channel uncertainty of the VU at time slot $t$. Therefore, the state space of the agent is ${s_{t}}=[\theta_{1}^{t},\cdots,\theta_{N}^{t},x_{1}^{t},y_{1}^{t},\cdots,x_{K}^{t},y_{K}^{t},{\gamma_{1}}(t-1),\cdots,{\gamma_{K}}(t-1)].$
@@
Initialize target networks by $\theta_{\mu^{\prime}}\leftarrow\theta_{\mu}$, $\theta_{Q^{\prime}}\leftarrow\theta_{Q}$;
@@ IV-B Concept of DDPG Algorithm
For high-dimensional state spaces, DDPG can be used to predict appropriate actions given the current state by deep neural networks that can learn complex features and patterns of the state. Therefore for the high-dimensional state and action space in this model, we use DDPG for training, and after obtaining the continuous action values, we can obtain the appropriate RIS phase-shift coefficients by scaling and rounding. In the DDPG algorithm, the state space and action space are usually represented by a deep neural network with two networks, actor and critic: the actor network is used to learn the strategy (i.e., action selection), and its goal is to maximize the estimation of the action value by the critic network, and the critic network is used to evaluate the value of the strategy, and learns to assess the evaluation by means of supervised learning merit of the actions output by the actor network. In addition, the DDPG algorithm uses a target actor network and a target critic network in order to ensure the stability of the algorithm. The network structure of these two networks is the same as that of the actor and critic networks, respectively.
Firstly, initialize the parameters $\theta_{\mu}$ and $\theta_{Q}$ of actor network $\mu(s;{\theta_{\mu}})$ and critic network $Q(s,a;{\theta_{Q}})$, and set the parameters $\theta_{\mu^{\prime}}$ and $\theta_{Q^{\prime}}$ of target actor network $\mu^{\prime}(s;\theta_{\mu^{\prime}})$ and target critic network $Q^{\prime}(s,a;\theta_{Q^{\prime}})$, respectively. Then, at each time slot $t$, action $a_{t}$ is selected based on the current strategy $\mu$ and exploration noise $\Delta n$, while the reward and next state are observed by executing the action. Next, the experience of each step is stored in the experience buffer for random sampling training. At each time slot $t$, when the buffer length is greater than $I$, a batch of data is sampled from the experience replay buffer $\mathcal{D}$ which is used to update the parameters of the critic network to minimize the loss function
@@
${\nabla_{{\theta_{\mu}}}}J\approx\frac{1}{{I}}\sum\limits_{i}{{\nabla_{a}}}Q(s,a;{\theta_{Q}}){|_{s={s_{i}},a=\mu({s_{i}})}}\nabla{\theta_{\mu}}\mu(s;{\theta_{\mu}}){|_{{s_{i}}}},$ (17)
@@ V Power Allocation Optimization By Modified MADDPG Algorithm
In this section, we will describe in detail the multi-agent environment and its associated states, actions and rewards, and finally, we will discuss the proposed MARL algorithm and its associated formulations. To be clear, our proposed framework for single-agent reinforcement learning and multi-agent reinforcement learning (SARL-MARL) is shown in Fig. 2, the relevant details are as follows.
@@ V-A Power Allocation Model for VEC
Algorithm 2 Modified MADDPG Power Allocation Algorithm with Joint RIS Phase-shift Model
In the previous section, the BS was used as an agent to interact with the environment and an optimal model for the RIS phase-shift was trained based on Algorithm 1. In this section, we consider the VU as an agent, apply the multi-agent approach to interact with the environment, and obtain the power allocation decision through the corresponding policy. In each training step of power allocation, we will use the corresponding phase-shift coefficients based on the previous RIS phase-shift model, and based on this, we will obtain the channel of the VU, and thus determine the part of the training state of the VU to proceed to the next training step. The relevant state space, action space and reward function in the multi-agent model are defined as follows:
State space: At time slot $t$, the state of each VU $k$ (agent $k$) consists of the following components: buffer length $q_{k}(t)$, the size of offloaded executed tasks $q_{k,o}(t)$, the size of locally executed tasks $q_{k,l}(t)$, and the offloaded and locally processed task overflows ${{q_{k,o}}(t)+{q_{k,l}}(t)-{q_{k}}(t)}$, the SNR of the $k$th VU at time slot $t$ ${\gamma_{k}}(t-1)$. Therefore, the state space of the VU $k$ at slot $t$ can be formulated as ${s_{k}(t)}=\left[{{q_{k}}(t),{q_{k,o}}(t),{q_{k,l}}(t),{q_{k,o}}(t)+{q_{k,l}}(t)-{q_{k}}(t),{\gamma_{k}}(t-1)}\right].$
Action space: as described above, the $k$th agent allocates the offloading power and the local power according to the corresponding policy, so that the action space of the $k$th agent at time slot $t$ is defined as ${a_{k}}(t)=\left[{{p_{k,o}}(t),{p_{k,l}}(t)}\right]$.
Reward function: in our proposed scheme, there are two aspects of rewards for each agent, one is the global reward reflecting the cooperation among agents, and the other is the local reward that is to help each agent to explore the optimal power allocation scheme. In DRL, rewards can be set flexibly, and a good reward can improve the performance of the system. Here, our goal is to optimize the agent $k$’s offloading and the local power level, as well as the corresponding buffer length. Thus the local and global rewards of the $k$th VU at time slot $t$ are
@@ V-B Multi-agent Deep Reinforcement Learning Algorithms
Based on DDPG algorithm, we propose a modified MADDPG algorithm that learns how to maximize both global and local rewards. It contains two main critics: a shared global critic and individual local critics. The global critic evaluates the states and actions of all agents collectively to optimize global rewards, while each agent has its own local critic that assesses rewards based on its specific local states and actions. Each agent operates with its own actor and critic networks for decision-making and action evaluation. Although agents can only observe their local environment, they can share experiences through a shared experience replay pool. During experience replay, agents store and sample from this shared pool of experiences, which enhances training efficiency and stability. This mechanism allows agents to learn from each other’s responses to the environment. The global reward signal promotes collaboration among agents by providing a unified reward feedback at each step, guiding the system towards optimized performance. In addition, we consider the impact of approximation errors in strategy and value updates on the global critic in MADDPG algorithm. Similar to [68], to improve the MADDPG algorithm, we employ twin-delay deterministic strategy gradient to replace the global critic.
Specifically, we consider a vehicular environment with $K$ vehicles (agents) and the policies for all agents are $\mathbf{\pi}=\{{\pi_{1}},{\pi_{2}},\cdots,{\pi_{K}}\}$. The $k$th agent’s strategy $\pi_{k}$, local critic ${\mathop{\rm Q}\nolimits}_{{\phi_{k}}}^{k}$ and twin global critic ${\mathop{\rm Q}\nolimits}_{\psi 1}^{{g_{1}}}$, ${\mathop{\rm Q}\nolimits}_{\psi 2}^{{g_{2}}}$ are parameterized by ${\theta_{k}}$, $\phi_{k}$, ${\psi_{1}}$ and $\psi_{2}$, respectively. For each agent, the modified policy gradient can be written as
@@
${y_{g}}={r_{g}}+\gamma{\mathop{\min}\limits_{j=1,2}}{\left.{Q_{\psi_{j}^{\prime}}^{{g_{j}}}({\bf{s^{\prime},a^{\prime}}})}\right|_{a_{k}^{\prime}=\pi_{k}^{\prime}(s_{k}^{\prime})}},$ (24)
where, ${\bf{\pi}}^{\prime}=\{\pi_{1}^{\prime},\pi_{2}^{\prime},\cdots,\pi_{K}^{\prime}\}$ is the target policy with parameter $\theta^{\prime}=\{\theta_{1}^{\prime},\theta_{2}^{\prime},\cdots,\theta_{K}^{\prime}\}$. Similarly, the local critic $Q_{{\phi_{k}}}^{k}$ of the $k$th agent is updated to
${L^{k}}({\phi_{k}})={{\rm E}_{{s_{k}},{a_{k}},{r_{k}},s_{k}^{\prime}}}\left[{{{(Q_{{\phi_{k}}}^{k}({s_{k}},{a_{k}})-y_{l}^{k})}^{2}}}\right],$ (25)
@@
The detailed MARL algorithm is described in the Algorithm 2. In addition, for the real-world implementation of Algorithm 2, it contains two aspects, on the one hand, for RIS phase-shift optimal coefficients, we need to obtain the vehicle’s position, RIS phase-shift coefficients, etc. in each training step, and obtain the relevant RIS phase-shift optimal model after training. On the other hand, since we use the MADDPG algorithm to optimize the offloading power and the local execution power of the VU, this enables the VU to complete the resource allocation locally, and we only need to use the RIS phase-shift optimal coefficients allocation results from the BS in each training step, which greatly reduces the communication cost.